# Node-Level Uncertainty Estimation in LLM-Generated SQL

### v1

@@ 1.  Introduction
However, Xiong et al., 2023 observed that LLMs can be overconfident in their responses, indicating that supervised uncertainty estimation could offer advantages when sufficient data is available. Since SQL generation is open-ended, we focus our approach on supervised node-level uncertainty quantification estimation. To that end, we begin by introducing the first semantically-aware labeling algorithm for correctness of nodes in pairs of generated and ground truth SQL – a priori an ambiguous task (See Section 3.1). See Appendix A Anonymous (2025) for how our design choices make the node-level correctness labeling agree with intuition in 13 cases we used while prototyping, while node-level hashing (the only alternative in related works) does not. We then create a training dataset with carefully designed, schema-informed, features and labels at the node level. (See Section 3.2.) Finally, we train a gradient-boosted tree model on this dataset to estimate node-level uncertainty. (See Section 4.) In particular, we interpret the term “uncertainty” as calibrated predictive probability of error.
@@
    In Section 3.1, given a generated SQL query and a ground truth SQL query, we give an algorithm for labeling what it means for a single node in the AST of a generated SQL to be correct (See Algorithm 1). See Appendix Anonymous (2025), which showcases how the design choices made the algorithm agree with intuition in 13 specific cases that we had used while prototyping.
@@ 3.1.  Ground Truth Generation
This traversal continues recursively through the structure of the generated AST, with alignment decisions based only on local subtree comparisons; the gold AST is not searched globally during this phase. As a result, some semantically correct nodes may remain labeled as errors due to greedy or structurally mismatched alignment decisions. To further improve the labeling heuristic, we do a final, context‑agnostic pass that globally re‑evaluates nodes still labeled as errors—trading structural precision for increased recall. The precise algorithm is spelled out in Algorithm 1; Appendix A Anonymous (2025) illustrates, across 13 cases, how these design choices align labels with intuitive judgments.
@@ 25      Relabel $n_{\text{gen}}$ as correct;
See Appendix A Anonymous (2025) to see how 5 out of the 13 unit tests we include there for node-level correctness labeling to agree with intuition would have failed if not for the post-processing.
@@ 5.  Bibliographical References
  • Anonymous (2025) Anonymous. 2025. Appendix for "node-level uncertainty estimation in llm-generated sql". https://anonymous.4open.science/r/LREC26-appendix-0452/lrec26_appendix.pdf. Anonymous submission, supplementary material.

### v2

@@ 1.  Introduction
However, Xiong et al., 2023 observed that LLMs can be overconfident in their responses, indicating that supervised uncertainty estimation could offer advantages when sufficient data is available. Since SQL generation is open-ended, we focus our approach on supervised node-level uncertainty quantification estimation. To that end, we begin by introducing the first semantically-aware labeling algorithm for correctness of nodes in pairs of generated and ground truth SQL – a priori an ambiguous task (See Section 3.1). See Appendix  A for how our design choices make the node-level correctness labeling agree with intuition in 13 cases we used while prototyping, while node-level hashing (the only alternative in related works) does not. We then create a training dataset with carefully designed, schema-informed, features and labels at the node level. (See Section 3.2.) Finally, we train a gradient-boosted tree model on this dataset to estimate node-level uncertainty. (See Section 4.) In particular, we interpret the term “uncertainty” as calibrated predictive probability of error.
@@
    In Section 3.1, given a generated SQL query and a ground truth SQL query, we give an algorithm for labeling what it means for a single node in the AST of a generated SQL to be correct (See Algorithm 1). See Appendix A, which showcases how the design choices made the algorithm agree with intuition in 13 specific cases that we had used while prototyping.
@@ 3.1.  Ground Truth Generation
This traversal continues recursively through the structure of the generated AST, with alignment decisions based only on local subtree comparisons; the gold AST is not searched globally during this phase. As a result, some semantically correct nodes may remain labeled as errors due to greedy or structurally mismatched alignment decisions. To further improve the labeling heuristic, we do a final, context‑agnostic pass that globally re‑evaluates nodes still labeled as errors—trading structural precision for increased recall. The precise algorithm is spelled out in Algorithm 1; Appendix  A illustrates, across 13 cases, how these design choices align labels with intuitive judgments.
@@ 25      Relabel $n_{\text{gen}}$ as correct;
See Appendix  A to see how 5 out of the 13 unit tests we include there for node-level correctness labeling to agree with intuition would have failed if not for the post-processing.